# A note for double Hölder regularity of the hydrodynamic pressure for weak solutions of Euler equations

### v3

@@ 4.2. Unbounded domains
In fact, we expect that this is valid when the last condition is replaced by the weaker condition: “${\rm II}$ has a uniform $C^{0}$-bounds over ${\partial\Omega}$”. In this case, one may say that the unbounded domain $\Omega$ has bounded $C^{2}$-geometry.
@@
This research was partially supported by NSFC under Grant No. 12331008, and the Shanghai Frontier Research Institute for Modern Analysis. The research of SL was also partially supported by NSFC under Grant Nos. 12201399 and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, National Key Research and Development Programs 2023YFA1010900 and 2024YFA1014900, and Shanghai Rising-Star Project. The research of YGW was also partially supported by NSFC under Grant Nos. 12171317, 12250710674 and 12161141004, and Shanghai Municipal Education Commission under Grant No. 2021-01-07-00-02-E00087.

### v4

@@ 4.2. Unbounded domains
In fact, we expect that this is valid when the last condition is replaced by the weaker condition: $\rm II$} has a uniform $C^{0}$-bounds over $\partial\Omega$}”. In this case, one may say that the unbounded domain $\Omega$ has bounded $C^{2}$-geometry.
@@
Funding Declaration. This research was partially supported by NSFC under Grant No. 12331008, and the Shanghai Frontier Research Institute for Modern Analysis. The research of SL was also partially supported by NSFC under Grant Nos. 12201399 and 12411530065, Young Elite Scientists Sponsorship Program by CAST 2023QNRC001, National Key Research and Development Programs 2023YFA1010900 and 2024YFA1014900, and Shanghai Rising-Star Project. The research of YGW was also partially supported by NSFC under Grant Nos. 12171317, 12250710674 and 12161141004, and Shanghai Municipal Education Commission under Grant No. 2021-01-07-00-02-E00087.
Author Contribution Declaration. SL and YGW wrote the main manuscript text and reviewed the manuscript.